OpenStudy (anonymous):

x^2-11x+18 Solution set

OpenStudy (saifoo.khan):

2,9

OpenStudy (anonymous):

is the answer 7.5 and 14.5?

OpenStudy (saifoo.khan):

No. 2,9

OpenStudy (anonymous):

please help with problem.

OpenStudy (saifoo.khan):

it's 2 and 9.

OpenStudy (anonymous):

don't know where i made mistake in quadratic equation?

OpenStudy (saifoo.khan):

u can simply factor it as well.

OpenStudy (saifoo.khan):

x^2-11x+18 x^2 - 9x -2x + 18 x(x-9) -2(x-9) = 0 (x-9)(x-2) = 0

OpenStudy (anonymous):

oh, right! much easier! thx

OpenStudy (saifoo.khan):

uw.